Session: Feb 2017 (session 1)
Wavelength: 0.7 cm


To give feedback on this experiment, click here

Experiment Feedback

Yebes (Angel Vicente at Sun Mar 12 04:17:11 2017) Success:
No problem

Effelsberg (Uwe at Mon Mar 13 09:50:52 2017) Success:
No problems, beside the broken LCP channel.

Onsala (Jun Yang at Wed Mar 15 13:32:44 2017) Success:
No known problems.